### v3

Abstract
The closure principle is a standard tool for achieving strong family-wise error rate (FWER) control in multiple testing problems. We develop an e-value-based closed testing framework that inherits nice properties of e-values, which are common in settings of sequential hypothesis testing or universal inference for irregular parametric models. We prove that e-value-based closed testing strongly controls the post-hoc FWER in the static setting, and has stronger anytime-valid and always-valid FWER-controlling properties in the sequential setting. Furthermore, we extend the celebrated graphical approach for FWER control (Bretz et al., 2009), using the weighted average of e-values for the local test, a strictly more powerful approach than weighted Bonferroni local tests with inverse e-values as p-values. In general, the computational cost for closed testing can be exponential in the number of hypotheses. Although the computational shortcuts for the p-value-based graphical approach are not applicable, we develop an efficient polynomial-time algorithm using dynamic programming for e-value-based graphical approaches with any directed acyclic graph, and tailored algorithms for the e-Holm procedure previously studied by Vovk and Wang (2021) and the e-Fallback procedure.
1 Introduction and Preliminaries
1.1 Family-wise error rate
In analyses that test multiple hypotheses, it is necessary to consider the problem of multiple testing, and control for some notion of false rejection. One such approach is to control the Family-wise Error Rate (FWER), the probability of any false rejection, at a given level $\alpha\in(0,1)$. In particular, given null hypotheses $H_{1},...,H_{n}$, with $\mathcal{H}_{0}$ denoting the subset of true nulls, an algorithm that produces the (random) rejection set $\mathcal{R}$ based on test statistics controls FWER at level $\alpha$ if
$\mathbb{P}(\mathcal{R}\cap\mathcal{H}_{0}\neq\emptyset)\leq\alpha.$
In the literature, this property is often referred to as strong FWER control, to distinguish from the weak form, which applies only when all null hypotheses are true. Since we will not address weak FWER control, except briefly in Section 2.2 to clarify existing work, we will adopt the simpler terminology throughout.
Controlling the FWER is crucial to maintain the reliability of statistical results, especially in medicine, economics, and online experimentation in tech firms. In medicine, clinical trials often compare the efficacy of several dosages or involve multiple endpoints, and FWER control ensures that no false positives undermine patient safety or lead to the adoption of ineffective therapies (Pocock et al., 1987; Bretz et al., 2009; Vickerstaff et al., 2019). In economics, researchers often test multiple hypotheses defined by multiple interventions, multiple demographic subgroups, and multiple policy outcomes at the same time; controlling FWER helps avoid erroneous conclusions that could misguide policy decisions (Romano and Wolf, 2005; List et al., 2019; Viviano et al., 2024). Similarly, in tech firms, A/B testing is used to compare multiple product versions or features under different metrics. Controlling FWER minimizes the chance of falsely identifying an intervention as superior, avoiding shipping ineffective features (Johari et al., 2022). Recently, FWER control has been applied to areas beyond traditional multiple testing problems, such as distribution-free risk control for black-box machine learning algorithms (Angelopoulos et al., 2021).
1.2 Closed testing
Given null hypotheses $H_{1},...,H_{n}$, a local test for index subset $I\subset[n]$, where $[n]$ is shorthand notation for the set $\{1,\ldots,n\}$, tests the hypothesis $H_{I}=\cap_{i\in I}H_{i}$ that all the nulls $H_{i}$ in $I$ are true. By convention, we call $H_{1},\ldots,H_{n}$ elementary hypotheses.
Given a family of valid local tests $\{\phi_{I}\}_{I\subset[n]}$, meaning $\mathbb{P}_{H_{I}}(\phi_{I}=1)\leq\alpha$ for all $I\subset[n]$, the “closure” of $\{\phi_{I}\}_{I\subset[n]}$ rejects $H_{i}$, the elementary hypothesis, if $\phi_{I}=1$ for all $I\ni i$. Any admissible procedure which controls FWER is a closed test (Sonnemann and Finner, 1988). The closure principle has also been extended to controlling more general error measures such as the false discovery rate (Xu et al., 2025; Goeman et al., 2025).
In the p-value literature, the most common local test is weighted Bonferroni, which given p-values $p_{1},...,p_{n}$, and a set of valid weights for any $I$, i.e. $\{w_{i}(I)\}$ with
$\sum_{i\in I}w_{i}(I)\leq 1,$ (1.1)
rejects $H_{I}$, or equivalently computes $\phi_{I}=1$ iff $\min_{i\in I}\frac{p_{i}}{w_{i}(I)}\leq\alpha$, which is valid as
$\mathbb{P}_{H_{I}}(\phi_{I}=1)=\mathbb{P}_{H_{I}}(\cup_{i\in I}\{p_{i}\leq\alpha w_{i}(I)\})\leq\sum_{i\in I}\mathbb{P}_{H_{I}}(p_{i}\leq\alpha w_{i}(I))\leq\sum_{i\in I}\alpha w_{i}(I)\leq\alpha$
by a union bound.
The challenge becomes simplifying the computation from the $12^{n}-$ 1 local tests to an efficient number. The literature includes many such procedures including Holm’s procedure, which uses the local test of unweighted Bonferroni with $w_{i}(I)=1/|I|$.
A larger class of local tests, in which Holm’s procedure belongs, is that of the graphical approach, introduced by Bretz et al. (2009). It is a generic framework that unifies most widely-used FWER controlling procedures and allows the researcher to encode logical relationships among hypotheses. We will revisit the local tests in detail in Section 5. Holm’s procedure is an example of the graphical approach, with the complete graph and equal weights. The Fallback procedure, introduced by Wiens and Dmitrienko (2005), is an example of the graphical approach with a chain graph. In general, any graph can be used for the purposes of the graphical approach, where its closure can be efficiently computed, with a simple greedy algorithm for determining the rejection set based on the p-values.
1.3 E-values
E-values are a recent alternative to p-values as a tool for hypothesis testing and quantifying evidence against the null (Ramdas and Wang, 2024; Ramdas et al., 2023; Waudby-Smith and Ramdas, 2024; Wang and Ramdas, 2022). For a null hypothesis $H_{0}$, an e-value for $H_{0}$ is a realization of an e-variable $e$ which has the property that $\mathbb{E}_{H_{0}}[e]\leq 1$. By Markov’s inequality, the test $\phi=1\{e\geq 1/\alpha\}$ is valid at level $\alpha$, since by Markov’s inequality $\mathbb{P}(e\geq 1/\alpha)\leq\alpha$.
E-values arise naturally in sequential settings, especially in always-valid inference, where the practitioner would like the flexibility of having valid inference at any stopping time. This framework is very relevant for the setting of online data collection in the tech sector, including the setting discussed by Johari et al. (2022). In these settings, the e-value is often defined through an e-process $(e_{it})_{t\geq 1}$, which is a stochastic process bounded by a non-negative martingale with marginal expectation $1$, known as a testing martingale (e.g. Ramdas et al., 2023).
E-values also prove to be useful in universal inference for irregular parametric models (Wasserman et al., 2020; Tse and Davison, 2022; Spector et al., 2023; Park et al., 2023). In general, e-values offer more robustness to get valid inference in the presence of various concerns about data-dependent experimental decisions, such as a post-hoc choice of $\alpha$ (Grünwald, 2024; Hemerik and Koning, 2024; Koning, 2024b).
1.4 Our contribution: e-value-based closed testing
In this paper, we first study closed testing procedures where each local hypothesis $H_{I}$ is associated with an e-value $e_{I}$. We prove that generic e-value-based closed tests control multiple stricter forms of Type-I error than the FWER. In particular, they achieve the post-hoc FWER control introduced by Koning (2024b), which allows researchers to choose the target FWER level after observing the data. In sequential settings where local e-values $e_{I}$ are given by e-processes (e.g. Ramdas et al., 2023), the resulting closed test guarantees the always-valid strong FWER control: with probability $1-\alpha$, no null hypothesis would ever be rejected if the test is applied at every time point.
Next, following the p-value-based closed testing literature, we investigate weighted e-Bonferroni local tests where the local e-value $e_{I}$ is given a weighted average of e-values corresponding to elementary hypotheses. We show that the weighted e-Bonferroni tests are provably more powerful than the weighted p-Bonferroni tests with the same weights and $1/e$ p-values, the only admissible derived p-values without distributional assumptions on the e-values (Wang, 2024). As a result, the closure of weighted e-Bonferroni tests is guaranteed to reject at least as many hypotheses as its p-value-based counterpart with $1/e$ p-values. We also examine the sequential settings where each elementary hypothesis $H_{i}$ is associated with an e-process $e_{it}$. In these settings, Ville’s inequality yields a more powerful p-value, known as the always-valid p-value, as the inverse of $\max_{t}e_{it}$, the global maximum of the e-process (Johari et al., 2022). While $\tilde{e}_{i}=\max_{t}e_{it}$ is not a valid e-value, we show that, if the e-processes associated with elementary hypotheses are independent, the e-value-based closed test applied to pseudo-e-values $\tilde{e}_{i}$ asymptotically controls FWER when the target level converges to zero and dominates the corresponding p-value-based closed test applied to always-valid p-values $1/\tilde{e}_{i}$.
Lastly, we develop efficient algorithms for e-graphical approaches, defined as the closure of weighted e-Bonferroni tests with the same weights used by p-graphical approaches and formally introduced in Section 2.4. The p-graphical approaches can be computed in polynomial time thanks to the consonance property of the weighted p-Bonferroni tests under the carefully constructed weights (Gabriel, 1969). The consonance property yields a sequential rejection algorithm with $O(n^{2})$ computational complexity. However, our e-value-based local tests are not consonant in general, as we discuss later in Section 5. While the standard sequential rejection algorithm does not work, we develop efficient polynomial time algorithms for e-graphical approaches with arbitrary directed acyclic graphs (DAG) using dynamic programming (Section 5). The computation complexity is $O(n|\mathcal{E}|)$ in general, where $|\mathcal{E}|$ is the number of edges, but usually smaller. We also improve the algorithm for two special cases: e-Holm (Section 3) and e-Fallback (Section 4). The e-Holm procedure was proposed in the prior work of Vovk and Wang (2021, 2023, 2024), though we derive derived a different representation based on the effective cutoff, computed in $O(n)$ time, which enables a direct comparison between e-Holm and p-Holm. Moreover, the previous work does not discuss other graphical approaches which pose substantially greater algorithmic challenges than e-Holm.
2 Theory of closed testing with e-values
2.1 E-value-based closed testing
In general, we define a e-value-based closed testing procedure as a collection of e-values $\{e_{I}\}_{I\subset[n]}$ where each $e_{I}$ is a valid e-value for the intersection hypothesis $H_{I}=\cap_{i\in I}H_{i}$. The procedure rejects $H_{i}$ iff $e_{I}\geq 1/\alpha$ for all supersets $I$ containing $i$.
For p-value closed testing, it is convenient to report an adjusted p-value $p_{i}^{*}$ for each hypothesis $H_{i}$ such that $H_{i}$ is rejected by the closed test iff $p_{i}^{*}\leq\alpha$. Analogously, we define the adjusted e-value
$e_{i}^{*}\triangleq\min_{I\ni i}e_{I}.$ (2.1)
Clearly, $H_{i}$ is rejected by the e-value-based closed test iff $e_{i}^{*}\geq 1/\alpha$.
These definitions yield an immediate yet crucial result that lays the foundation of all error-controlling properties of e-value-based closed tests.
Lemma 2.1.
Let $e_{i}^{*}$ be adjusted e-values defined in (2.5). Then
$\max_{i\in\mathcal{H}_{0}}e_{i}^{*}\leq e_{\mathcal{H}_{0}}.$
Since $e_{\mathcal{H}_{0}}$ is a valid e-value, we can easily prove the FWER controlling property.
Proposition 2.2.
The e-value-based closed test controls the FWER:
$\mathrm{FWER}=\mathbb{P}\left(\max_{i\in\mathcal{H}_{0}}e_{i}^{*}\geq\frac{1}{\alpha}\right)\leq\mathbb{P}\left(e_{\mathcal{H}_{0}}\geq\frac{1}{\alpha}\right)\leq\alpha.$
In the following two subsections, we leverage Lemma 2.1 to prove stronger error-controlling properties of e-value-based closed testing.
2.2 Post-hoc FWER control
Grünwald (2024) formalizes the notion of roving $\alpha$s, or post-hoc choice of $\alpha$ at which to control the error rate. They prove that if using an e-value $e$ to test a single hypothesis, one can interpret the data at any (potentially data-dependent) level since for any $\hat{\alpha}$ which is a function of the data,
$\mathbb{E}\left[\frac{\mathbf{1}\{e\geq\frac{1}{\hat{\alpha}}\}}{\hat{\alpha}}\right]\leq\mathbb{E}[e]\leq 1,$
Koning (2024b) generalizes the post-hoc type-I error control to multiple testing. In particular, the Appendix C.2 of Koning (2024b) introduces the notion of post-hoc weak FWER control in terms of controlling the probability of a false rejection when all hypotheses of interest are nulls, i.e., $\mathcal{H}_{0}=\{1,\ldots,n\}$.
We prove that the post-hoc strong FWER control can be achieved by e-value-based closed testing.
Theorem 2.3.
For closed testing adjusted e-values $\{e_{i}^{*}\}_{i=1}^{n}$ and set of null hypotheses $\mathcal{H}_{0}$, for any data-dependent function of the data $\hat{\alpha}$,
$\mathbb{E}\left[\frac{\mathbf{1}\{\max_{i\in\mathcal{H}_{0}}e_{i}^{*}\geq\frac{1}{\hat{\alpha}}\}}{\hat{\alpha}}\right]\leq 1$
In general, a p-value defined by $\mathbb{P}[p\leq\alpha]\leq\alpha$ is not post-hoc valid, so we would expect a p-value-based closed test to not control post-hoc FWER. As a concrete example, consider a collection of independent p-values with $n_{0}$ uniform null p-values $p_{i}:i\in\mathcal{H}_{0}$. Then the p-Holm procedure rejects $H_{i}$ iff its adjusted p-value $p_{i}^{*}\triangleq\max_{I\ni i}\min_{j\in I}|I|p_{j}$ is below $\alpha$. If we choose $\hat{\alpha}=\min_{i}p_{i}^{*}$, the minimal level at which at least one hypothesis is rejected,
$\mathbb{E}\left[\frac{\mathbf{1}\{\min_{i\in\mathcal{H}_{0}}p_{i}^{*}\leq\hat{\alpha}\}}{\hat{\alpha}}\right]$ $=\mathbb{E}\left[\frac{1}{\min_{i\in\mathcal{H}_{0}}p_{i}^{*}}\right]\geq\mathbb{E}\left[\frac{1}{n_{0}\min_{j\in\mathcal{H}_{0}}p_{j}}\right]$
$\geq\frac{1}{n_{0}}\frac{1}{1/(n_{0}+1)}=1+\frac{1}{n_{0}}>1.$
We obtain the first inequality since by the closure principle, $\min_{i\in\mathcal{H}_{0}}p_{j}^{*}\leq p_{\mathcal{H}_{0}}=n_{0}\min_{j\in\mathcal{H}_{0}}p_{j}$, and the second inequality by Jensen’s inequality since the minimum of $n_{0}$ uniforms has mean $1/(n_{0}+1)$.
2.3 Stronger FWER control in sequential settings
In this section, we consider e-value-based closed testing in the sequential setting. Let $\mathcal{F}_{t}$ denote the $\sigma$-field generated by data up to time $t$ collected for all hypotheses. For each intersection hypothesis $H_{I}$, we define an e-process $e_{It}$ with respect to this filtration $\mathcal{F}_{t}$. Wang et al. (2025) refer to e-processes of this kind as global e-processes, in contrast to local e-processes that are valid with respect to their own filtrations. Then we can also define the adjusted e-processes $e^{*}_{it}$ as the adjusted e-values at time $t$. By Lemma 2.1, we have that for each $t$,
$\max_{i\in\mathcal{H}_{0}}e^{*}_{it}\leq e_{\mathcal{H}_{0},t}.$ (2.2)
Since $e_{\mathcal{H}_{0},t}$ is an e-process, the optional stopping theorem implies that $e_{\mathcal{H}_{0},\tau}$ is a valid e-value for any $\mathcal{F}_{t}$-stopping time $\tau$. Using the same argument as in Theorem 2.3, we prove the post-hoc-anytime-validity.
Theorem 2.4.
For any $\mathcal{F}_{t}$-stopping time $\tau$ and $\mathcal{F}_{\tau}$ measurable level $\hat{\alpha}$,
$\mathbb{E}\left[\frac{\mathbf{1}\{\max_{i\in\mathcal{H}_{0}}e_{i\tau}^{*}\geq\frac{1}{\hat{\alpha}}\}}{\hat{\alpha}}\right]\leq 1$
While anytime-validity enables more flexible timing for decisions, it forbids making decisions based on retrospective inspection. By Ville’s inequality, we can prove the always-validity of e-value-based closed testing with a fixed target level.
Theorem 2.5.
For any fixed level $\alpha\in(0,1)$,
$\mathbb{P}\left(\max_{i\in\mathcal{H}_{0}}\sup_{t}e_{it}^{*}\geq\frac{1}{\alpha}\right)\leq\alpha$
As a result of Theorem 2.5, the strong FWER of any null adjusted e-value ever being rejected is controlled by $\alpha$. This means that we can make rejections based on the supremum over $t$ over the adjusted e-values, even if the supremum was in the past. This property is stronger than anytime-validity; see Ramdas et al. (2020) for a detailed comparison between two notions of sequential validity.
Theorem 2.5 no longer holds for general data-driven $\alpha$. Following the counterexample from the proof of Proposition 3.1 of Tavyrikov et al. (2025), we construct two e-processes $(e_{1t},e_{2t})$ where $e_{10}=e_{20}=1$ and $e_{it}=\prod_{s=1}^{t}e_{s}^{i}$, where
$(e_{s}^{1},e_{s}^{2})=\begin{cases}(1/2,1/2)&\text{w.p. }1/3\\ (1/2,2)&\text{w.p. }1/3\\ (2,1/2)&\text{w.p. }1/3.\end{cases}$
When run with the running maxima $\tilde{e}_{it}=\max_{0\leq s\leq t}e_{is}$, e-Holm, the closure of e-Bonferroni with equal weights, makes a rejection if $\tilde{e}_{1t}+\tilde{e}_{2t}\geq 2/\alpha$ and either $\tilde{e}_{1t}\geq 1/\alpha$ or $\tilde{e}_{2t}\geq 1/\alpha$. In this way, for any $t>0$, we can set $\hat{\alpha}_{t}=2/(\tilde{e}_{1t}+\tilde{e}_{2t})$, which is guaranteed to be below $1$ as $\tilde{e}_{it}\geq e_{i0}=1$, and we are guaranteed that at time $t$, at least one hypothesis, $i_{t}=\arg\max_{i\in\{1,2\}}\tilde{e}_{it}$, will be rejected at level $\hat{\alpha}_{t}$. In this case, the family-wise scaled risk is
$\mathbb{E}\left[\frac{\mathbf{1}\{\text{reject a hypothesis at level }\hat{\alpha}_{t}\}}{\hat{\alpha}_{t}}\right]=\mathbb{E}\left[\frac{\tilde{e}_{1t}+\tilde{e}_{2t}}{2}\right]=\mathbb{E}[\tilde{e}_{1t}]$
by symmetry. It is clear that $\mathbb{E}[\tilde{e}_{1t}]>\mathbb{E}[e_{1t}]=1$. For example, taking $t=1$, this procedure will have inflated post-hoc risk, since $\tilde{e}_{1t}\geq 1$, and with probability 1/3, $\tilde{e}_{11}=2$, so $\mathbb{E}[\tilde{e}_{11}]\geq\frac{4}{3}>1$.
Nonetheless, following Tavyrikov et al. (2025), we can apply adjusters to running maxima $\tilde{e}_{It}=\max_{s\leq t}e_{Is}$ for all intersection hypotheses to convert them to e-processes with respect to the same filtration (e.g. Shafer et al., 2011; Koolen and Vovk, 2014; Choe and Ramdas, 2024). The adjusted closed test achieves post-hoc always-valid FWER control.
Theorem 2.6.
Let $A$ be any increasing function satisfying
[…]
Then applying the adjuster $A$ to the e-value-based closed testing adjusted maximums of e-processes $\tilde{e}_{i}^{*}$, yielding $A(\tilde{e}_{i}^{*})$, controls FWER for any data-driven level $\hat{\alpha}$ measurable in $\bigcup_{t\geq 1}\mathcal{F}_{t}$, i.e.,
$\mathbb{E}\left[\frac{\mathbf{1}\{\max_{i\in\mathcal{H}_{0}}A(\tilde{e}_{i}^{*})\geq\frac{1}{\hat{\alpha}}\}}{\hat{\alpha}}\right]\leq 1.$
According to Choe and Ramdas (2024) and Wang et al. (2025), $\{A(e_{It}):I\subset[n]\}$ are global e-processes even if $\{e_{It}:I\subset[n]\}$ are local e-processes with respect to a coarser filtration than $\mathcal{F}_{t}$. In such cases, Theorem 2.4 continues to hold.
In Theorem 2.6, there are three operations at play, namely applying the maximum over time, adjusting via closed testing, and calibrating with the adjuster $A$. We apply $A$ last, but it is also valid to apply the adjuster $A$ first then compute the closure, as $A(e_{it})$ is a valid global e-process. In practice, most popular adjusters are concave, including $A(e)=\sqrt{e}-1$ and $A(e)=\frac{e-1-\log e}{(\log e)^{2}}$ (Tavyrikov et al., 2025). For weighted e-Bonferroni local tests, the Jensen’s inequality implies $A(\tilde{e}_{I})=A(\sum_{i\in I}w_{i}(I)\tilde{e}_{i})\geq\sum_{i\in I}w_{i}(I)A(\tilde{e}_{i})$ for each subset $I$, so applying the adjuster after closure leads to higher power than the reverse.
2.4 Weighted e-Bonferroni local test
To obtain a valid local test with these e-values, one approach is to convert them into p-values using e-to-p calibrators (Vovk and Wang, 2021). Their Proposition 2.2 shows that the unique admissible converted p-value is the inverse e-value truncated at $1$, i.e., $\min\{1/e_{i},1\}$, which is further formalized in Wang (2024). We will ignore the truncation without loss of generality, as none of the procedures discussed in this paper reject p-values greater than or equal to $1$. Then the weighted Bonferroni test rejects $H_{I}$ iff
$\min_{i\in I}\frac{1}{w_{i}(I)e_{i}}\leq\alpha\Longleftrightarrow\max_{i\in I}w_{i}(I)e_{i}\geq\frac{1}{\alpha}.$ (2.3)
While (2.3) is a valid test, it does not fully unleash the potential of e-values. An appealing feature of e-values is that they combine easily; the weighted average of evalues is an e-value, as noted by Vovk and Wang (2021). It is also possible to combine independent e-values by multiplication, which Fischer and Ramdas (2024) use with closed testing, but we do not make any independence between e-values assumptions.
This suggests an alternative local test, which we call weighted e-Bonferroni due to its analogousness to weighted p-Bonferroni, that rejects $H_{I}$ iff
$e_{I}\triangleq\sum_{i\in I}w_{i}(I)e_{i}\geq\frac{1}{\alpha}.$ (2.4)
The test is valid as $e_{I}$ is a valid e-value.
When the weights $w_{i}(I)$ are chosen according to a p-graphical approach, we call the closed test an e-graphical approach. The nomenclature extends to specific graphical approaches, such as e-Holm and e-Fallback.
Comparing (2.3) with (2.4), it is evident that the weighted e-Bonferroni test is more powerful in the sense that it rejects $H_{I}$ whenever the weighted p-Bonferroni test (2.3) does. As a consequence, the e-value-based closed test is at least as powerful as its inverse-e p-value counterpart.
The weighted e-Bonferroni local test, with weights $w_{i}(I)$, leads to the following optimization to compute adjusted e-values:
$e_{i}^{*}=\min_{I\ni i}\left\{\sum_{i\in I}w_{i}(I)e_{i}\right\}.$ (2.5)
Vovk and Wang (2021, 2023) proposed an efficient algorithm for e-Holm where $w_{i}(I)=1/|I|$ in our notation. We develop dynamic programming-based polynomial-time algorithms for more general e-graphical approaches.
2.5 More powerful e-value-based closed testing with e-processes
In the sequential setting discussed in Section 2.3, Johari et al. (2022) introduced the always-valid p-value, defined as $1/\tilde{e}_{i}$ where $\tilde{e}_{i}=\sup_{t\geq 0}e_{it}$. In general, $\tilde{e}_{i}$ is not a valid e-value (e.g. Tavyrikov et al., 2025) – this does not contradict the admissibility of the inverse-e e-to-p calibrator (Vovk and Wang, 2021) because e-processes impose restrictions on the e-values.
When using weighted e-Bonferroni local tests, the e-value-based closed testing procedure described in Section 2.3 does not dominate closed testing based on always-valid p-values, as the latter are more powerful than inverse-e p-values. Nevertheless, we can still apply the e-value-based closed testing to the pseudo-e-values $\tilde{e}_{i}$, following Theorem 2.6, and in the same spirit of applying the Benjanimi-Hochberg procedure to running maximums of e-processes (Tavyrikov et al., 2025). Specifically, we define
$\tilde{e}_{I}=\sum_{i\in I}w_{i}(I)\tilde{e}_{i},$
and reject all $H_{i}$ with $\tilde{e}_{i}^{*}\geq 1/\alpha$, where $\tilde{e}_{i}^{*}$ is computed according to (2.1).
We show that, when the e-processes are independent, this procedure asymptotically controls the FWER as $\alpha\rightarrow 0$. The properties of e-value-based procedures with $\alpha\rightarrow 0$ have also been studied in other contexts (Koning, 2024a; Koning and van Meer, 2025).
Theorem 2.7.
Assume that $\{(e_{it})_{t\geq 1}:i\in\mathcal{H}_{0}\}$ are independent e-processes with respect to the filtration $\mathcal{F}_{t}$ defined at the beginning of Section 2.3. Then the closed testing adjusted pseudo-e-values asymptotically control FWER for small $\alpha$:
$\lim_{\alpha\rightarrow 0}\frac{\mathbb{P}(\max_{i}\tilde{e}_{i}^{*}\geq\frac{1}{\alpha})}{\alpha}\leq 1.$ (2.6)
3 e-Holm
3.1 A simple expression of the rejection rule
The Holm’s procedure is the closure of a set of unweighted p-Bonferroni local tests with weight $w_{i}(I)=1/|I|$ for every subset $I\subset[n]$ and $i\in I$. Following the description in Section 2.4, the e-Holm procedure closes the unweighted e-Bonferroni local tests:
$H_{I}\text{ rejected iff }\frac{1}{|I|}\sum_{i\in I}e_{i}\geq\frac{1}{\alpha}.$
By definition of the closure principle, the e-Holm procedure rejects $H_{i}$ iff
$\sum_{j\in I}e_{j}\geq\frac{|I|}{\alpha}\text{ for all }I\ni i.$ (3.1)
Surprisingly, the rejection rule boils down to a simple thresholding rule with an easy-to-compute critical value, which can be computed in $O(n)$ time.
Theorem 3.1.
The e-Holm procedure rejects $H_{i}$ iff
$e_{i}\geq\frac{1}{\alpha}+C,\quad\text{where }C=\sum_{j=1}^{n}\max\left\{\frac{1}{\alpha}-e_{j},0\right\}.$ (3.2)
To compute the adjusted e-values, we re-derived a $O(n\log n)$ algorithm of Vovk and Wang (2021, 2023), where the $\log n$ term is from sorting. For completeness, we restate and explain the intuition behind this algorithm in Appendix D.
Notably, the threshold rejection procedure (3.2) can be computed without sorting the e-values, so Theorem 3.1 does provide a runtime improvement over the adjusted e-values algorithm if the goal is to recover the rejection set for a fixed $\alpha$.
3.2 Comparing e-Holm with p-Holm
The p-Holm procedure with inverse-e p-values rejects $H_{(i)}$ iff
$\frac{1}{e_{(j)}}\leq\frac{\alpha}{n-j+1},\quad\text{for any }j\leq i.$
The discussion in Section 2.4 already implies that (3.2) is less stringent. To bring more insight, we facilitate the comparison by considering an extreme case where the non-nulls are strong and $e_{i}>1/\alpha$ almost surely (or with extremely high probability). In order for p-Holm to make any rejection, the largest e-value $e_{(1)}$ needs to exceed $n/\alpha$. By contrast, the threshold $1/\alpha+C$ in e-Holm (3.2) is bounded by $(|\mathcal{H}_{0}|+1)/\alpha$ from above, which can be much smaller than $n/\alpha$. Therefore, e-Holm can adapt to the number of nulls while p-Holm cannot. This is analogous to adaptive false discovery rate control (Storey et al., 2004; Benjamini et al., 2006).
4 e-Fallback
4.1 A naive dynamic programming algorithm
We start by presenting an suboptimal yet simple algorithm to illustrate how we apply dynamic programming to simplify the computation. The Fallback procedure is defined by a fixed sequence of hypotheses, $H_{1}\rightarrow H_{2}\rightarrow\cdots\rightarrow H_{n}$ with associated e-values, and initial $\alpha$ budget $\{\alpha_{i}\}_{i=1}^{n}$ with $\sum_{i=1}^{n}\alpha_{i}=\alpha$. This is a special case of the later discussed graphical case with a chain graph. For $|I|=k$ with $I=\{i_{1},...,i_{k}\}$ and $i_{1}<i_{2}<\cdots<i_{k}$, the local test rejects $H_{I}$ iff
$e_{I}\triangleq\sum_{\ell=1}^{k}w_{i_{\ell}}(I)e_{i_{\ell}}\geq\frac{1}{\alpha},\quad\text{where }w_{i_{\ell}}(I)=\sum_{i=i_{\ell-1}+1}^{i_{\ell}}\frac{\alpha_{i}}{\alpha}.$ (4.1)
Above, we take $i_{0}=0$ for notational simplicity. This formula corresponds to redistributing $\alpha$ budget from excluded indices to the nearest subsequent included index.
We first prove that the adjusted e-values only depend on subsets with $i$ being the largest element. This result extends to the general DAGs (see Theorem 5.4 in Section 5) and turns out to be a crucial property that enables an efficient dynamic programming approach.
Lemma 4.1.
For each $i\in[n]$,
$e_{i}^{*}=\max_{I\ni i,i=\max_{k\in I}k}e_{I}.$
Based on this property, we give Algorithm 1, a dynamic programming algorithm with $O(n^{2})$ runtime to compute the minimum over $I\ni i$ for each $i\in[n]$.
Input: Vectors $\{e_{i}\}_{i=1}^{n},\{\alpha_{i}\}_{i=1}^{n}$;
Initialize: $m_{0}=0$;
Iterate: $m_{i}=\min_{0\leq j<i}\{m_{j}+e_{i}\sum_{k=j+1}^{i}\alpha_{k}\}$;
Output: $\{m_{i}\}_{i=1}^{n}$.
Algorithm 1 e-Fallback
We now prove that our computation of $m_{i}$ gives the adjusted e-values.
Theorem 4.2.
The output of Algorithm 1 has the property that $m_{i}=\alpha e_{i}^{*}$ for all $i\in[n]$.
This dynamic programming approach performs a minimization over at most $n$ values for each of $n$ indices so has $O(n^{2})$ runtime. However, as we will see the next two subsections, we can achieve an $O(n)$ runtime with some algorithmic shortcuts.
4.2 Speedup via a reverse search algorithm
In this subsection, we show that there is an algorithm to perform the optimization of e-Fallback without searching all $i$ previous hypotheses each time we compute $m_{i}$. This algorithm will serve as a stepping stone to the graphical approaches with general DAGs in Section 5, as well as the tailored algorithm for e-Fallback in Section 4.3.
For any $i\in I\subset[i]$, instead of viewing $e_{I}$ as a weighted average of e-values, we can interpret it as a weighted average of initial $\alpha$-budgets $\alpha_{1},\ldots,\alpha_{i}$, each appearing exactly once in the expression. For any such subset $I=\{i_{1},\ldots,i_{k}\}$ with $0=i_{0}<i_{1}<i_{2}<\ldots<i_{k}=i$, it divides $[i]$ into $k$ pieces and assigns the weight $e_{i_{\ell}}$ to the $\ell$-th piece $\{\alpha_{i_{\ell-1}+1},\ldots,\alpha_{i_{\ell}}\}$ in the formula of $e_{I}$. Adding an element $j\in(i_{\ell-1},i_{\ell})$ in $I$ amounts to cut the $\ell$-th chunk into two and reassigns the weight $e_{j}$ to $\alpha$-budgets in $\{i_{\ell-1}+1,\ldots,j\}$. To achieve the minimum, one should never add $j$ with $e_{j}>e_{i_{\ell}}$. This observation implies a recursive structure of the optimal subset for each hypothesis.
Theorem 4.3.
For each $i\in[n]$, let $j(i)=\max\{j<i:e_{j}\leq e_{i}\}$ if the set is nonempty and $j(i)=0$ otherwise. Define $I_{i}^{*}$ iteratively as follows:
$I_{0}^{*}=\emptyset,\quad I_{i}^{*}=I_{j(i)}^{*}\cup\{i\}.$
Then
$e_{i}^{*}=e_{I_{i}^{*}}=\frac{1}{\alpha}\left(\sum_{j=j(i)+1}^{i}\alpha_{j}\right)e_{i}+e_{j(i)}^{*}.$
Theorem 4.3 leads to the following reverse search algorithm. Again $\{m_{i}/\alpha\}_{i=1}^{n}$ is our set of adjusted e-values.
Input: Vectors $\{e_{i}\}_{i=1}^{n},\{\alpha_{i}\}_{i=1}^{n}$;
Initialize: $m_{i}=0$ for all $i=0...n$;
Iterate: For $i=1$ to $n$:
Do: Set $j=i-1$ and repeat until $e_{j}\leq e_{i}$:
$m_{i}=m_{i}+\alpha_{j}e_{i}$, $j=j-1$
$m_{i}=m_{i}+m_{j}$.
Output: $\{m_{i}\}_{i=1}^{n}$.
Algorithm 2 Reverse Search for e-Fallback
This algorithm gives a significant shortcut when the previous smaller e-value $e_{j}$ is close to $e_{i}$. If, for example, we assume that the e-values are continuous and exchangeable, then the expected runtime is $O(n)$. We state this as Theorem C.1, which we prove in Section C. However, in the worst case when the e-values are decreasing, the runtime, based on the total number of back searches, is still $O(n^{2})$. In the next subsection, we will utilize a stack to recycle intermediate quantities more efficiently and reduce the computational complexity to $O(n)$.
4.3 A tailored algorithm for e-Fallback
Inspecting the construction of the optimal subset $I_{i}^{*}$ in Theorem 4.3, we observe that iff
$j\in I_{i}^{*}\Longleftrightarrow e_{j}=\min\{e_{j},\ldots,e_{i}\},$ (4.2)
that is, $e_{j}$ is a cumulative minimum defined in a backward manner. $I_{i}^{*}$ is the lower envelope of the first $i$ e-values. We illustrate it in Figure 1. This implies the following iterative construction of $I_{i}^{*}$ in a similar spirit to Lemma D.1 for e-Holm.
Input: Vectors $\{e_{i}\}_{i=1}^{n},\{\alpha_{i}\}_{i=1}^{n}$;
Initialize: $m_{i}=0$ for all $i=0...n$, stack $s=\{\}$;
Iterate: For $i=1$ to $n$:
Do: Set $\alpha_{i}^{*}=\alpha_{i}$:
Pop $(j,a_{j}^{*})$ from $s$ and add $\alpha_{i}^{*}=\alpha_{i}^{*}+\alpha_{j}^{*}$ until $e_{j}\leq e_{i}$ or $s=\{\}$
If stop because $e_{j}\leq e_{i}$: $m_{i}:=\alpha_{i}^{*}e_{i}+m_{j}$, push $(j,\alpha_{j}^{*})$ to $s$
If stop because $s=\{\}$: $m_{i}:=\alpha_{i}^{*}e_{i}$
push $(i,\alpha_{i}^{*})$ to $s$
Output: $\{m_{i}\}_{i=1}^{n}$.
Algorithm 3 Stack Search for e-Fallback
Lemma 4.4.
Suppose $I_{i}^{*}$ defined in Theorem 4.3 is $\{k_{1}(i),\ldots,k_{m_{i}}(i)\}$ where $m_{i}=|I_{i}^{*}|$. Then $e_{k_{1}(i)}\leq\ldots\leq e_{k_{m_{i}}(i)}$ and
$I_{i}^{*}=\{i\}\cup I_{i-1}^{*}\setminus\{j\in I_{i-1}^{*}:e_{j}>e_{i}\}.$
The construction in Lemma 4.4 can be implemented by a stack, as detailed in Algorithm 3.
Figure 1: Visualization of stack e-Fallback Algorithm 3. For $H_{8}$, $I_{8}^{*}$ includes all four orange squares. At step $i=9$, the 7th and 8th squares are kicked out because their e-values are above $e_{9}$ and hence are no longer cumulative minimums. The 4th and 5th squares are kept because $e_{5}$ is the most recent square whose e-value is below $e_{9}$. Going into step 10, $e_{9}$ will be added to the stack.
As desired, Algorithm 3 has runtime $O(n)$. To see this, we break down the computation into push and pop operations. Since at most two elements are pushed at the end of each iteration, the total cost of push operations is $O(n)$. Using the representation (4.2), all but the last popped elements at any iteration will not be added back to the stack again. The cost of last pop adds up to $O(n)$, while the cost of all other pops combined cannot exceed the number of hypotheses $n$. Therefore, the overall computational complexity is $O(n)$.
5 General approach for DAGs
5.1 Revisiting the p-graphical approach
A p-graphical approach takes as input an initial $\alpha$-budget $\alpha_{i}$ for $H_{i}$, with $\sum_{i}\alpha_{i}=\alpha$ and a weighted directed graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, with $\mathcal{V}=\{H_{1},\ldots,H_{n}\}$. Each vertex encodes a hypothesis and the weight $q_{jk}$ between $H_{j}$ and $H_{k}$ is nonnegative and satisfying $\sum_{k}q_{jk}\leq 1$, where $q_{jk}>0$ if and only if $(j,k)\in\mathcal{E}$. To define the weights $w_{i}(I)$ for the local test $H_{I}$, expand the graph with an additional vertex, denoted by $H_{n+1}$, with graph weight $q_{j(n+1)}=1-\sum_{k}q_{jk}$ and $q_{(n+1)j}=I(j=n+1)$. Consider the random walk $Z_{0},Z_{1},\ldots$ on the expanded graph with
$\mathbb{P}(Z_{0}=H_{i})=\frac{\alpha_{i}}{\alpha},\quad\mathbb{P}(Z_{t+1}=H_{k}\mid Z_{t}=H_{j})=q_{jk}.$
Then, the weight $w_{i}(I)$ is chosen as
$w_{i}(I)=\mathbb{P}(Z_{t_{I}}=H_{i})$ (5.1)
where $t_{I}=\min\{t:Z_{t}\in\{H_{i}:i\in I\}\}$.
As mentioned in Section 1.4, the closure of the above local tests can be computed in polynomial time due to the consonance property, which states that, for any $I\subset[n]$, $H_{I}$ is rejected if there exists $i\in I$ such that $H_{J}$ is rejected for any $i\in J\subset I$. To see this, $w_{i}(I)\leq w_{i}(J)$ for any $J\subset I$. When $H_{I}$ is rejected, there must be an $i\in I$ such that $p_{i}\leq\alpha w_{i}(I)$. Then, for any $i\in J\subset I$, $p_{i}\leq\alpha w_{i}(J)$ and hence $H_{J}$ is rejected. The graph is allowed to be cyclical.
However, the weighted e-Bonferroni local tests are not consonant. For example, in the context of e-Holm, with three hypotheses, desired level $\alpha=0.05$, and e-values $(e_{1},e_{2},e_{3})=(25,25,10)$, the subset $\{1,2,3\}$ is rejected as the average e-value is $e_{\{1,2,3\}}=20\geq\frac{1}{\alpha}$. However, there is no $i$ that satisfies the definition of the consonance property; in particular, $e_{\{1,3\}}=e_{\{2,3\}}=17.5<\frac{1}{\alpha}$.
5.2 Dynamic programming for e-DAG graphical approaches
For a DAG, the hitting probabilities can be simplified using the unidirectional flow of the graph. To facilitate the computation, for any subset $I$ define the set of paths from $j\notin I$ to $i\in I$ through the complement of $I$:
$\mathcal{P}_{j,i}^{(I)}=\{(i_{0},...,i_{k}):k\geq 0,i_{0}=j,i_{k}=i,i_{\ell}\notin I\,\,\forall\ell\in[k-1],q_{i_{\ell-1},i_{\ell}}>0\,\,\forall\ell\in[k]\}.$ (5.2)
Note the positivity of a graph weight ($q_{i_{\ell-1},i_{\ell}}>0$) indicates the presence of an edge. Implicit in this definition is that $i\in I,j\notin I$. Also define $k(p):=k$ the highest index of a path $p=(i_{0},...,i_{k})\in\mathcal{P}_{j,i}^{(I)}$. Then,
$w_{i}(I)=\frac{1}{\alpha}\left(\alpha_{i}+\sum_{j\notin I}\alpha_{j}\sum_{p\in\mathcal{P}_{j,i}^{(I)}}\prod_{\ell=1}^{k(p)}q_{i_{\ell-1},i_{\ell}}\right).$ (5.3)
We may think of the local test weight corresponding to redistributing the weight ($\alpha_{j}/\alpha$) from each excluded vertex $j$ proportionally according to the graph structure. With the weight $w_{i}(I)$ defined in (5.3), the local test for e-DAG graphical approaches rejects $H_{I}$ iff
$e_{I}\triangleq\sum_{i\in I}w_{i}(I)e_{i}\geq\frac{1}{\alpha}.$ (5.4)
For a general DAG, we notice a local phenomenon that we claim will extend to an algorithm that computes the global minimum for any index $i$. We first establish an equivalence between the local test e-value $e_{I}$ as a weighting of e-values $e_{i}$ for $i\in I$, and a weighting of alpha values $\alpha_{j}$ for $j\in[n]$.
Definition 5.1.
For any subset $I\subset[n]$ and hypothesis index $j\in[n]$, define
$e_{j}^{(I)}\triangleq\begin{cases}\sum_{i\in I}e_{i}\left(\sum_{p\in\mathcal{P}_{j,i}^{(I)}}\prod_{\ell=1}^{k(p)}q_{i_{\ell-1},i_{\ell}}\right)&j\notin I\\ e_{j}&j\in I\end{cases}$ (5.5)
When no path exists, we take $e_{j}^{(I)}=0$.
Intuitively, we give this definition to give structure to the way in which an excluded vertex $i\notin I$ will have its alpha-budget $\alpha_{j}$ redistributed to the e-values in $I$ via a reweighting based on the graph structure. Formally, we prove this relationship in Lemma 5.2.
Lemma 5.2.
For $e_{j}^{(I)}$ as defined in Definition 5.1 and subset $I\subset[n]$, we have
$e_{I}=\sum_{i\in I}w_{i}(I)e_{i}=\frac{1}{\alpha}\sum_{j=1}^{n}\alpha_{j}e_{j}^{(I)}$ (5.6)
The result of Lemma 5.2 is useful as it shows that if we can find a subset $I\ni i$ that minimizes $e_{j}^{(I)}$ for all $j$, then we have found the minimizing subset, $\mathrm{argmin}_{I\ni i}e_{I}$.
To facilitate the upcoming proof of our algorithm’s validity, we also show the following Lemma, which says that we may recursively define $e_{j}^{(I)}$ by its children’s values.
Lemma 5.3.
For $j\notin I$,
$e_{j}^{(I)}=\sum_{(j,k)\in E}q_{j,k}e_{k}^{(I)}$ (5.7)
We assume that the nodes are already topologically ordered with respect to the input graph, which is a reasonable constraint given in practice a practitioner will be designing the graph themselves so will be able to design the data collection to have the nodes ordered properly.
We give the extension of the reverse search Fallback algorithm from Section 4.2 to the general DAG in Algorithm 4. The intuition is that we can look solely at the ancestor graph, and search backward to decide whether or not to include each node, thereby reconstructing the minimizing subset for each node $i$.
Input: Vectors $\{e_{i}\}_{i=1}^{n},\{\alpha_{i}\}_{i=1}^{n}$, transition matrix $(q_{ij})\in\mathbb{R}^{n\times n}$, where $[n]$ is a
            topological ordering with respect to the graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$;
Iterate: For $i\in[n]$:
Do: Initialize e-assignments $e_{j}^{(i)}=0\,\,(j\in[n])$, $m_{i}=0$, compute $A_{i}$ the set of
       ancestors of $i$, including $i$;
Initialize $e_{i}^{(i)}=e_{i}$, $m_{i}=\alpha_{i}e_{i}$:
For $i\neq j\in A_{i}$ decreasing $e_{j}^{(i)}=\min\left(e_{j},\sum_{(j,k)\in E,k\in A_{i}}q_{jk}e_{k}^{(i)}\right)$ and set $m_{i}=m_{i}+\alpha_{j}e_{j}^{(i)}$.
Output: $\{m_{i}\}_{i=1}^{n}$
Algorithm 4 e-DAG Graphical Approach
Theorem 5.4.
In the context of Algorithm 4, for every index $i$ and subset $I\ni i$, $e_{j}^{(I)}\geq e_{j}^{(i)}$. Moreover, $e_{j}^{(i)}=e_{j}^{(I_{i}^{*})}$ for some $I_{i}^{*}\ni i$, which is therefore the minimizing subset. As a result, $e_{i}^{*}=m_{i}/\alpha$.
In Algorithm 4, for each node $i$, we recompute the e-value assignments for all $j\leq i$, each of which is a comparison of $e_{j}$ to a weighting of its children’s assignments. The total computational complexity is $O\left(\sum_{i}N_{i}\right)$ where $N_{i}$ is the number of the edges in the ancestor graph $A_{i}$. This yields a worst-case complexity bound $O(n|\mathcal{E}|)=O(n^{3})$ because $N_{i}\leq|\mathcal{E}|=O(n^{2})$. This worst-case bound can be improved for special graphs. For example, when $\mathcal{G}$ is a tree graph, $N_{i}$ is simply the depth of node $H_{i}$. For a balanced binary tree, $\max_{i}N_{i}=O(\log n)$, and thus the complexity is $O(n\log n)<\!\!<O(n|\mathcal{E}|)$. It may be possible to further improve the complexity by re-using intermediate calculations like Algorithm 3. For example, in the tree setting, each reverse search is on a chain, so each reverse search can use the shortcut from Algorithm 2, and in particular we believe one can use a depth-first search along with a stack, following Algorithm 3, to achieve an $O(n)$ runtime. We leave further developments for future research.
6 Simulation results
6.1 Sequential setting
In this section we develop a framework for comparing the power of our e-value procedures to that of the corresponding p-value procedures. Without further restrictions, Vovk and Wang (2021) suggest that the only valid p-values are given by inverse e-values. Thus, our e-graphical approaches are strictly more powerful than the corresponding p-graphical approaches. To make the comparison more informative, we work in the sequential testing setting, where we can derive a valid p-value that is more powerful than the inverse e-value (Johari et al., 2022).
We present results for power in terms of how often the e-procedure improves on the p-procedures, as well as by what percentage in terms of ratio of stopping times. Both give useful information for gauging how meaningful the potential improvement of e-Holm is in the sequential setting. We examine the Holm’s procedure described in Section 3 and the graphical procedure described in Section 5, specifically for a useful structuring of hypotheses in a factorial design.
For both of our simulation settings, we will be operating in a setting of a continuous stream of Gaussian data $Y_{it}\overset{i.i.d.}{\sim}\mathcal{N}(\mu_{i},1)$, with known variance 1. For arm $i$, i.i.d. data comes in at times $t=1,2,...$, with the same mean $\mu_{i}$. In Section 6.2, we will be testing each arm separately, while in Section 6.3 we will be testing differences of means of arms. In either case, the way we will construct test martingales is through the sequential probability ratio test (SPRT) with known alternative, a la Johari et al. (2022). In particular, let $\mathcal{F}_{t}$ denote the $\sigma$-field generated by $\{(X_{1j},\ldots,X_{nj}):j\leq t\}$. For any given predictable sequence $\hat{\mu}_{s}\in\mathcal{F}_{s-1}$, let, we let
$e_{it}=\prod_{s=1}^{t}\frac{\phi(Y_{is}-\hat{\mu}_{i})}{\phi(Y_{is})}=\prod_{s=1}^{t}\exp(\hat{\mu}_{i}Y_{is}-0.5\hat{\mu}_{i}^{2}),$ (6.1)
where $\phi$ is the pdf (likelihood) of the Gaussian distribution. Clearly, $e_{it}$ is a martingale with $\mathbb{E}[e_{it}]=1$. By optional stopping theorem, for any stopping time $\tau$ with respect to the filtration $\mathcal{F}_{t}$, $\mathbb{E}[e_{i\tau}]=1$. As a result, $e_{i\tau}$ is a valid e-value and, hence, $1/e_{i\tau}$ is a valid p-value. By Ville’s inequality (Ville, 1939), $1/\max_{t\leq\tau}e_{it}$ is also a valid p-value. Johari et al. (2022) call it an always-valid p-value. We will perform e-graphical approaches on $e_{i\tau}$ and p-graphical approaches on $1/e_{i\tau}$ and $1/\max_{t\leq\tau}e_{it}$, where $T$ is an appropriately defined stopping time, adapted to $\mathcal{F}_{t}$.
The multiple testing procedures, when applied to the e-processes, provide always-valid FWER control, similar to reasoning from Johari et al. (2022). Because $\{e_{it}\}_{t\geq 1}$ is a valid martingale adapted to $\mathcal{F}_{t}$ for all $i\in[n]$, the stopped processes $\{e_{i\tau}\}_{i\in[n]}$ are e-values and the closure principle will provide FWER control. Any stopping criterion that uses the adjusted e-values from current or previous times $t$ is a function of the processes, and so is still adapted to the full filtration $\mathcal{F}_{t}$. However, if some processes are adapted to sub-filtrations, the validity may be an issue since the stopped processes may not all be e-values. Choe and Ramdas (2024) discuss examples of encountering and correcting for this issue.
6.2 e-Holm simulations
Our setting is 20 hypotheses $H_{i}:\mu_{i}=0$, for i.i.d. data $Y_{it}\sim\mathcal{N}(\mu_{i},1)$, where the known alternative for each is $\mu_{alt}$, which we vary on $\{0.5,1,1.5,2\}$. For our simulation, there are 5 true alternatives. We define the test martingale by (6.1) with $\hat{\mu}_{i}=\mu_{alt}$ and our data collection stopping time as the time at which at least one hypothesis can be rejected, after accounting for multiplicity (via each procedure). In terms of adjusted e-values $\{e_{it}^{*}\}_{i=1}^{n}$, this stopping time is $\tau_{e}=\min\{t:\max_{i\in[n]}e_{it}^{*}\geq 1/\alpha\}$, and we define $\tau_{ep}$ and $\tau_{p}$ analagously for the 1/e p-values and always-valid p-values.
For each value of $\mu_{alt}$, we run 1000 Monte Carlo simulations, where we save the stopping times $\tau_{e},\tau_{ep},\tau_{p}$, all computed with the same sequence of simulated data. For practical purposes there is a maximum number of iterations of 2000, though no run reaches this threshold. Because we stop once the most significant hypothesis is rejected, its test martingale is at its maximum, and thus $\tau_{e}\leq\tau_{p}$ always, and $\tau_{p}=\tau_{ep}$. This follows from the discussion of the comparison between e-Bonferroni and p-Bonferroni in Section 1.3. Thus, we only compare $\tau_{e}$ and $\tau_{p}$, and compute for $m=1000$ iterations and empirical distribution $\mathbb{P}_{m}$:
  • Comparison (probability of improving on p-Holm): $\mathbb{P}_{m}(\tau_{e}<\tau_{p})$
  • Ratio (conditional on there being a difference): $\mathbb{E}_{m}[\tau_{e}/\tau_{p}|\tau_{e}\neq\tau_{p}]$
For the former, shown in the left panel of Figure 2, the e-Holm comparison is always non-negative, the greater the better. For the latter, shown in the right panel of Figure 2, the e-Holm ratio is always no greater than 1, the lower the better. Both metrics are necessary to understand the efficiency of e-Holm, as a small ratio is meaningless with a small comparison.
Figure 2: Left: Empirical comparison from $m=1000$ of the probability of e-Holm improving on p-Holm. Right: Empirical ratio of stopping times, when they differ, between e-Holm and p-Holm using 1/e p-values (red) and always-valid p-values (blue). Both comparisons are for a stopping time of the first rejection out of 20 hypotheses.
The probability comparison in Figure 2 shows that the biggest improvement in likelihood of beating p-Holm comes for smaller $\mu_{alt}$, which makes intuitive sense as the stopping times are larger, leaving more margin for improvement. This reasoning also applies to the results in the right panel of Figure 2, which show that the lower the stopping time the lower the potential the ratio has to be, as low as 0.6 on average for $\tau_{e}\neq\tau_{p}$ when $\mu_{alt}=2$. This intuitive rationalization aside, it is promising that all offer improvements of at least $0.05$ in terms of $\mathbb{P}_{m}(\tau_{e}<\tau_{p})$ and about $10\%$ for $\mathbb{E}_{m}[\tau_{e}/\tau_{p}|\tau_{e}\neq\tau_{p}]$.
6.3 e-DAG simulations
We continue our simulation study with a similar framework under a different model, this time with a structure to our hypotheses in comparison to the agnostic treatment of hypotheses in Section 6.2. We use the example of a factorial design, which is a useful framework for testing not just primary nodes, but also secondary, tertiary, etc. nodes. For example, one might be interested in the effect of a drug and the effect of gender on outcomes as primary node, but also the interaction effect of the drug for a particular gender, a secondary node.
To illustrate this framework, consider a simple three-factor model, say for A/B testing in a tech application, with three proposed treatments. The model is that data $Y_{i}$ is drawn with treatments $X_{i}\in\{0,1\}^{3}$ according to
$Y_{i}=\beta_{0}$ $+\beta_{1}X_{i1}+\beta_{2}X_{i2}+\beta_{3}X_{i3}$
$+\beta_{12}X_{i1}X_{i2}+\beta_{23}X_{i2}X_{i3}+\beta_{13}X_{i1}X_{i3}$ (6.2)
$+\beta_{123}X_{i1}X_{i2}X_{i3}+\epsilon_{i},$
where $\epsilon_{i}\sim\mathcal{N}(0,\sigma^{2})$ is an i.i.d. Gaussian error term. In practice one might use a more complicated model but for simplicity we take this linear model with fixed effects and i.i.d. errors with known variance (which we take to be 1). Here the hypotheses $H_{j}:\beta_{j}=0$ correspond to primary nodes, $H_{jk}:\beta_{jk}=0$ correspond to secondary interaction nulls, and $H_{123}:\beta_{123}=0$ corresponds to the tertiary interaction null.
In an experiment, the practitioner likely cares mostly about the primary hypotheses, but would like to test the secondary (and higher order) hypotheses as well. An approach is to use the graphical procedure to define a graph in which each primary hypothesis is a root, whose children are the secondary hypotheses containing the primary hypothesis. The children of $H_{1}$ are $H_{12}$ and $H_{13}$ in our three-way example, which is visualized in Figure 3.
Figure 3: The transition graph for the graphical approach in a factorial design with three primary nodes.
In the context of the graphical approach, it is most common to take the transition probabilities to be uniform, and the starting $\alpha$ budget to be $\alpha/p$ for each of the $p$ primary endpoints, in our case $\alpha/3$ for each of $H_{1},H_{2},H_{3}$. The result of this choice is that, in terms of rejecting any primary hypothesis, we use Bonferroni to determine if any node can be rejected. Once one of these nodes can be rejected, its budget transfers to the corresponding secondary hypotheses, which might be rejected, though they require a higher level of significance. The ancestor graph of $H_{i}$ is just $H_{i}$, so we can only reject $H_{i}$ if its standalone test martingale can reject at level $\alpha/3$.
The upside is that rejecting, for example, $H_{13}$, becomes potentially easier because we can use e-Bonferroni as the local test, as if $H_{1}$ is rejected, we now can boost out adjusted e-value for $H_{13}$ via the value of $H_{3}$, if it remains unrejected. However, it is unclear a priori whether this, or the fact that the always-valid p-value gets to use the max value of the test martingale, will win out in various settings.
In order to investigate this trade-off, we perform a similar simulation study to Section 6.2, with our three-way factorial design model. It is the same in that we have $m=1000$ iterations, and we vary $\mu_{alt}\in\{0.5,1,1.5,2\}$. Our stopping time is the rejection of $H_{13}$, where $\beta_{1},\beta_{3},\beta_{13}$ are the only nonzero parameters. We fix $\beta_{1}=\beta_{3}=0.5$, with this alternative being known, and $\beta_{13}$ being unknown.
We consider a process in which at time $t$, we receive $82^{3}=$ 8 measurements corresponding to draws of our model with each of $X\in\{0,1\}^{3}$, which we will notate using binary. We can then isolate each coefficient (plus mean zero noise) from the data by taking combinations of our time $t$ data. For example, $Y_{t100}-Y_{t000}\sim\mathcal{N}(\beta_{1},2)$, and similar for the other primary endpoints. Then to isolate $\beta_{13}$, we can take
$Y_{t101}-Y_{t100}-Y_{t001}+Y_{t000}\sim\mathcal{N}(\beta_{13},4).$ (6.3)
Because we are adding and subtracting more datapoints, the variance increases. This procedure generalizes via a pattern that follows the inclusion-exclusion principle. Then for each of these test statistics, we can construct a test martingale via a SPRT for an alternative mean, which for us is known except for $H_{13}$, for which we estimate $\hat{\mu}_{t}$ based on the sample average before time $t$.
With these test martingales defined, we define our stopping time to be the time that $H_{13}$ is rejected, after accounting for multiplicity via e-DAG/ep-DAG/p-DAG, i.e. $\tau_{e}=\min\{t:e_{t}^{13*}\geq 1/\alpha\}$, for $e_{t}^{13*}$ the adjusted e-value for $H_{13}$, and analagously defined for $\tau_{p}$ and $\tau_{ep}$. Each of these procedures only depends on the test martingales for $H_{1},H_{3},H_{13}$.
As in Section 6.2, we record the probability of $\tau_{e}$ improving on $\tau_{p}$, but now it could be worse, and $\tau_{p},\tau_{ep}$ could differ, so we record three comparisons: $\mathbb{P}_{m}(\tau_{e}<\tau_{p}),\mathbb{P}_{m}(\tau_{e}>\tau_{p})$, and $\mathbb{P}_{m}(\tau_{e}<\tau_{ep})$. Since e-values have other documented pros besides our procedure as outlined in Section 1.3, and demonstrated in our e-value-based closed testing results in Section 2, knowing the latter is relevant. We also record the ratio conditional on a difference, both for $\tau_{p}$ and $\tau_{ep}$, because they might differ. In notation, these quantities are $\mathbb{E}_{m}[\tau_{e}/\tau_{p}|\tau_{e}\neq\tau_{p}]$ and $\mathbb{E}_{m}[\tau_{e}/\tau_{ep}\tau_{e}\neq\tau_{ep}]$.
We do two simulations. The first is for primary budget, meaning each of $H_{1},H_{2},H_{3}$ is assigned budget $\alpha/3$, and no budget for the secondary and tertiary hypotheses. We also perform the same analysis for equal budget among all seven hypotheses, which allows $H_{13}$, the subject of the simulation study, to be rejected even if $H_{1}$ and $H_{3}$ are both unrejected. This might be desirable if the secondary hypothesis is of interest separately from the primary ones. The probability comparisons are given in Figure 4 and the stopping time ratios are given in Figure 5.
In general, power tends to more favorable for e-DAG for higher values of $\mu_{alt}$, with the only favorable ratio occurring for $\mu_{alt}=2$. However, the results show that the gains over ep-DAG are consistent and substantial, with settings in which e-DAG is able to improve upon p-DAG.
Just like the simulations show, the equal budget experiment generally follow similar trends to plots for primary budget, but with improvements across the board. Most notably, the probability of p-DAG beating e-DAG becomes very low for $\mu_{alt}=2$, with a corresponding ratio of less than 0.95. Both simulation settings (primary and equal budget for factorial design) show that e-DAG can consistently be a notable improvement on ep-DAG, while having mixed results compared to p-DAG, albeit with many promising settings.
Figure 4: Empirical probability from $m=1000$ iterations of the e-DAG stopping time being less than ep-DAG (red), less than p-DAG (green), or greater than p-DAG (blue). The stopping time is when we reject a specific secondary hypothesis in a factorial design graphical model. With primary budget and equal budget respectively.
Figure 5: Empirical ratio from $m=1000$ iterations of the stopping times, when they differ, from Figure 4 for e-DAG against p-DAG (red) and ep-DAG (blue). With primary budget and equal budget respectively.
7 Conclusion and discussion
In this paper, we formalized the framework of e-value-based closed testing, in which each local test in the closure principle for FWER control is based on an e-value. We showed that e-value-based closed testing provides the ability to select the level of control $\alpha$ post-hoc, both in the traditional and anytime-valid settings. We also show e-value-based closed testing applied to any global e-processes provides always-valid FWER control for fixed $\alpha$ and that applied to the running maximums and calibrated with an adjuster achieves both post-hoc and always validity. Furthermore, when $\alpha\rightarrow 0$, the closure of weighted e-Bonferroni local tests applied to pseudo-e-values as inverse running maximums of e-processes provides strong FWER control.
A large class of closed tests are based on weighted p-Bonferroni based on a graph structure. We extend this graphical approach to weighted e-Bonferroni by developing efficient algorithms to compute the rejection set for adjusted e-values. For the procedure based on unweighted Bonferroni, we build on the e-Holm algorithm of Vovk and Wang (2021, 2023) to get a simple rejection threshold, and quantify the power increase compared to p-Holm with inverse-e p-values. For the graphical approach on directed acyclic graphs, we develop an efficient algorithm to compute the e-DAG closed test, and refine the algorithm to an $O(n)$ algorithm for the e-Fallback procedure which is based on a chain graph.
7.1 Improved thresholds for local tests
In earlier sections, we focused on the standard e-value test that uses a threshold of $1/\alpha$ for each local hypothesis. Recent studies have demonstrated that this threshold can be uniformly improved. Specifically, Ramdas and Manole (2023) and Koning (2024a) show that $1/\alpha$ can be replaced by $U/\alpha$ where $U$ is uniformly distributed on $[0,1]$ and its randomness is exogenous. For closed testing, we can define the local test as $e_{It}\geq U_{I}/\alpha$ for a collection of uniform random variables $\{U_{I}:I\subset[n]\}$ that are independent of the e-values. By Lemma 2.1, the FWER is bounded by
$\mathbb{P}\left(e_{\mathcal{H}_{0}}\geq\frac{U_{\mathcal{H}_{0}}}{\alpha}\right)=\mathbb{E}[\alpha e_{\mathcal{H}_{0}}]\leq\alpha.$
As a result, the randomized e-value-based closed testing controls FWER for any arbitrary dependence structure among the uniforms. In the sequential setting, if we use a time-invariant uniform random variable for each local test, we can similarly prove the anytime-valid FWER control. However, it remains unclear whether other properties discussed in Section 2, such as the post-hoc or always-valid FWER control, continue to hold.
With additional distributional assumptions on the e-values, Blier-Wong and Wang (2024) show that the threshold could be improved in a deterministic fashion. For example, if the e-value has a decreasing density, its $(1-\alpha)$-th quantile is bounded by $2/\alpha$ (Theorem 2 (i) of Blier-Wong and Wang (2024)). Applying this result to our setting, if $e_{1},\ldots,e_{n}$ for elementary hypotheses are independent with decreasing densities, any weighted average of them still has a decreasing density. As a consequence, the closure of any collection of weighted e-Bonferroni test at level $2\alpha$ controls FWER at $\alpha$.
7.2 FWER control under logical constraints
In previous sections, we focus on the case with logically independent hypotheses – that is, any combination of true and false nulls is considered possible. However, in many real-world applications, the hypotheses are logically related, meaning that some combinations of true/false status are ruled out a priori (Shaffer, 1986; Goeman and Solari, 2010; Loper et al., 2022). Logical constraints are deterministic rules that restrict the allowable configurations of null and alternative hypotheses. For example, under a multi-factor model like (6.3), the hereditary principle is often invoked as a structural prior that interactions are unlikely to be present unless supported by main effects. More formally, the strong/weak hereditary principle states that $H_{ij}$ must be null if $H_{i}$ and/or $H_{j}$ are correct nulls. Under the logical constraints, the intersection hypothesis $H_{I}$ is equivalent to $H_{J(I)}$ where $J(I)\subset I$ is the minimal subset of $I$ such that no hypothesis in $J(I)$ is implied by any hypothesis in $I$. For example, under the strong hereditary principle, $J(I)=\{\{1\},\{2\}\}$ when $I=\{\{1\},\{2\},\{1,2\},\{1,3\},\{1,2,3\}\}$. The logical constraints hence lead to a reduction of the number of hypotheses – the adjusted e-value can be defined as $e_{i}^{*}=\min_{I\ni i}e_{J(I)}$, which is at least as large as the adjusted e-value in the absence of logical constraints. All theoretical properties discussed in Section 2 remain to hold for this procedure. Nonetheless, the dynamic programming-based algorithm for e-graphical approaches needs modification. We leave this for future research.
7.3 Post-hoc false discovery proportion inference with e-values
The influential work by Goeman and Solari (2011) proposes a generic approach to construct a simultaneous confidence bound on the false discovery proportion (FDP) based on closed testing. Specifically, for each given subset $\mathcal{R}\subset[n]$, let
$\ell_{1-\alpha}(\mathcal{R})=\min\{|\mathcal{R}\setminus J|:\phi_{I}=0\,\,\forall I\supset J\}.$
Then $\ell_{1-\alpha}(\mathcal{R})$ is a uniform lower confidence bound on the number of true discoveries in $\mathcal{R}$:
$\mathbb{P}\left(|\mathcal{R}\setminus\mathcal{H}_{0}|\geq\ell_{1-\alpha}(\mathcal{R}),\,\,\forall\mathcal{R}\subset[n]\right)\geq 1-\alpha.$
For p-value-based closed testing, the consonance property leads to fast algorithms (Goeman and Solari, 2011). Vovk and Wang (2023, 2024) develop an efficient algorithm for the e-Holm procedure, leveraging the symmetry of the unweighted e-Bonferroni procedure. For general e-graphical approaches, we can compute $\ell_{1-\alpha}(\mathcal{R})$ by brute force when $n$ is small. For large $n$, the computation becomes more involved due to the asymmetry and we leave this for future research.
Appendix A Proofs
A.1 Proofs of results in Section 2
Proof of Theorem 2.3.
We can upper bound this quantity by replacing $\hat{\alpha}$ with $\alpha$ and taking the supremum, and moreover, by Lemma 2.1, we can write out
$\mathbb{E}\left[\frac{\mathbf{1}\{\max_{i\in\mathcal{H}_{0}}e_{i}^{*}\geq\frac{1}{\hat{\alpha}}\}}{\hat{\alpha}}\right]\leq\mathbb{E}\left[\sup_{\alpha}\frac{\mathbf{1}\{\max_{i\in\mathcal{H}_{0}}e_{i}^{*}\geq\frac{1}{\alpha}\}}{\alpha}\right]\leq\mathbb{E}\left[\sup_{\alpha}\frac{\mathbf{1}\{e_{\mathcal{H}_{0}}\geq\frac{1}{\alpha}\}}{\alpha}\right].$
By Grünwald [2024], this right hand quantity is bounded by 1 since $e_{\mathcal{H}_{0}}$ is a valid e-value and thus post-hoc valid with scaled error $\leq 1$ in expectation. However, we can reconstruct a simple proof of this fact by seeing that $\mathbf{1}\{e_{\mathcal{H}_{0}}\geq\frac{1}{\alpha}\}/\alpha$ is maximized when the numerator is 1 and $\alpha$ is as large as possible, which happens when $\alpha=1/e_{\mathcal{H}_{0}}$. This computes the upper bound as
$\mathbb{E}\left[\sup_{\alpha}\frac{\mathbf{1}\{e_{\mathcal{H}_{0}}\geq\frac{1}{\alpha}\}}{\alpha}\right]=\mathbb{E}\left[\frac{1}{1/e_{\mathcal{H}_{0}}}\right]=\mathbb{E}[e_{\mathcal{H}_{0}}]\leq 1.$
Proof of Theorem 2.4.
The proof follows from (2.2) and the properties of $e_{\mathcal{H}_{0},\tau}$ as an e-value, so we may bound
$\mathbb{E}\left[\frac{\mathbf{1}\{\max_{i\in\mathcal{H}_{0}}e_{i\tau}^{*}\geq\frac{1}{\hat{\alpha}}\}}{\hat{\alpha}}\right]$ $\leq\mathbb{E}\left[\frac{\mathbf{1}\{e_{\mathcal{H}_{0},\tau}\geq\frac{1}{\hat{\alpha}}\}}{\hat{\alpha}}\right]\leq\mathbb{E}\left[\sup_{\alpha}\frac{\mathbf{1}\{e_{\mathcal{H}_{0},\tau}\geq\frac{1}{\alpha}\}}{\alpha}\right]$
$=\mathbb{E}\left[\frac{1}{1/e_{\mathcal{H}_{0},\tau}}\right]=\mathbb{E}\left[e_{\mathcal{H}_{0},\tau}\right]\leq 1$
Proof of Theorem 2.5.
We first swap the supremum over $t$ and maximum over $i$, then appeal to the stochastic dominance (2.2).
$\mathbb{P}\left(\max_{i\in\mathcal{H}_{0}}\sup_{t}e_{it}^{*}\geq\frac{1}{\alpha}\right)=\mathbb{P}\left(\sup_{t}\max_{i\in\mathcal{H}_{0}}e_{it}^{*}\geq\frac{1}{\alpha}\right)\leq\mathbb{P}\left(\sup_{t}e_{\mathcal{H}_{0},t}\geq\frac{1}{\alpha}\right)\leq\alpha,$
where the last inequality follows from Ville’s inequality.
Proof of Theorem 2.6.
By Ville’s inequality, $1/\sup_{t}e_{It}$ is a valid p-value for the intersection hypothesis $H_{I}$. Let $B(x)=A(1/x)$. The condition on $A$ is equivalent to
[…]
Thus, $A(\sup_{t}e_{It})=B(1/\sup_{t}e_{It})$ is a valid e-value. Applying the same argument as in Theorem 2.3, the proof is completed. ∎
Proof of Theorem 2.7.
We generate independent uniform random variables $U_{i},i\in\mathcal{H}_{0}$. Since $\tilde{e}_{i}$ are independent, $\tilde{e}_{\mathcal{H}_{0}}$ stochastically dominates their weighted sum:
$\tilde{e}_{\mathcal{H}_{0}}=\sum_{i\in\mathcal{H}_{0}}w_{i}(\mathcal{H}_{0})\tilde{e}_{i}\succeq\sum_{i\in\mathcal{H}_{0}}w_{i}(\mathcal{H}_{0})U_{i}^{-1}.$
By Theorem 2 (ii) of Chen et al. [2023],
$\lim_{\alpha\rightarrow 0}\frac{\mathbb{P}\left(\sum_{i\in\mathcal{H}_{0}}w_{i}(\mathcal{H}_{0})U_{i}^{-1}\geq 1/\alpha\right)}{\alpha}=1.$
The proof is then completed by Lemma 2.1 which implies $\max_{i\in\mathcal{H}_{0}}\tilde{e}_{i}^{*}\leq\tilde{e}_{\mathcal{H}_{0}}$.
Proof of Theorem 3.1.
We can rewrite the rejection rule (3.1) as
$e_{i}\geq\frac{1}{\alpha}+\max_{I\ni i}\sum_{j\in I\setminus\{i\}}\left(\frac{1}{\alpha}-e_{j}\right).$ (A.1)
Taking $I=\{i\}$ implies that $H_{i}$ would not be rejected when $e_{i}<1/\alpha$. Thus we assume $e_{i}\geq 1/\alpha$ throughout the rest of the proof. It is clear that the RHS of (A.1) is bounded by $1/\alpha+C$ from above. On the other hand, it equals $1/\alpha+C$ if $I=\{i\}\cup\{j:e_{j}<1/\alpha\}$. The proof is then completed. ∎
A.2 Proofs of results in Section 4
Proof of Lemma 4.1.
For any $I\ni i$, $J=I\cap[i]$ has $\alpha e_{I}\geq\alpha e_{J}$, since the $\alpha$-budget for $j>i$ gets assigned nowhere. Specifically, taking $i=i_{\ell^{\prime}}$ without loss of generality,
$\alpha e_{I}-\alpha e_{J}=\sum_{\ell=\ell^{\prime}+1}^{k}\left(\sum_{i=i_{\ell-1}+1}^{i_{\ell}}\alpha_{i}\right)e_{i_{\ell}}\geq 0.$
Proof of Theorem 4.2.
By Lemma 4.1, we only need to show that $m_{i}=\min_{i\in I\subset[i]}\alpha e_{I}$. For each $i=i_{k}$, we consider the various options for $i_{k-1}$, which can be any element of $[i-1]$ or $k=1$ so $I=\{i\}$. Thus we can rewrite
$\alpha e_{i}^{*}=\min_{i\in I\subset[i]}\alpha e_{I}=\min_{0\leq j<i}\left[\min_{j\in J\subset[j]}\alpha e_{J}+e_{i}\sum_{k=j+1}^{i}\alpha_{k}\right],$
where the minimum for $j=0$ is 0 by convention. Then we use induction, with our base case of $m_{1}=\alpha_{1}e_{1}$, where we assume that $\min_{j\in J\subset[j]}\alpha e_{J}=m_{j}$ for all $j\leq i-1$, in which case we get that $\min_{i\in I\subset[i]}\alpha e_{I}=m_{i}$, as intended.
Proof of Theorem 4.3.
We prove the result by induction. The case for $i=1$ is evident by Lemma 4.1. Suppose the result holds for $11,\ldots,i-$ 1. If $j(i)=0$, $e_{i}>e_{j}$ for all $j<i$ and $I_{i}^{*}=\{i\}$. Then, for any $i\in I\subset[i]$,
$e_{I}\geq\left(\sum_{j=1}^{i}\alpha_{j}\right)e_{i}=e_{I_{i}^{*}}.$
For the rest of the proof, we assume $j(i)>0$.
Let $I_{i}=\{i_{1},\ldots,i_{k}\}$, with $i_{1}<i_{2}<\ldots<i_{k}=i$, be any subset such that $e_{i}^{*}=e_{I_{i}}$. We first prove that there exists $I_{i}^{\prime}$ such that $e_{i}^{*}=e_{I_{i}^{\prime}}$ and $j(i)$ is the second largest element. Assume that $j(i)\in(i_{\ell-1},i_{\ell}]$ for some $\ell\leq k$, let $I_{i}^{\prime}=\{i_{1},\ldots,i_{\ell-1},j(i),i\}$. Then
$\alpha e_{I_{i}}-\alpha e_{I_{i}^{\prime}}=\sum_{r=\ell}^{k}\left(\sum_{j=i_{r-1}+1}^{i_{r}}\alpha_{j}\right)e_{i_{r}}-\left(\sum_{j=i_{\ell-1}+1}^{j(i)}\alpha_{j}\right)e_{j(i)}-\left(\sum_{j=j(i)+1}^{i}\alpha_{j}\right)e_{i}$
$=\left(\sum_{j=i_{\ell-1}+1}^{j(i)}\alpha_{j}\right)(e_{i_{\ell}}-e_{j(i)})+\left(\sum_{j=j(i)+1}^{i_{\ell}}\alpha_{j}\right)(e_{i_{\ell}}-e_{i})+\sum_{r=\ell+1}^{k}\left(\sum_{j=i_{r-1}+1}^{i_{r}}\alpha_{j}\right)(e_{i_{r}}-e_{i}).$
By definition of $j(i)$, $e_{i_{r}}>e_{i}\geq e_{j(i)}$ for any $r\geq\ell$. Thus, $e_{I_{i}}\leq e_{I_{i}^{\prime}}$. Since $I_{i}$ is a minimizer, $I_{i}^{\prime}$ must be so as well.
Since we can assume $j(i)$ is the second largest element in $I_{i}$, $e_{i}^{*}$ can be written as
$e_{i}^{*}=\frac{1}{\alpha}\left(\sum_{j=j(i)+1}^{i}\alpha_{j}\right)e_{i}+\min_{j(i)\in J\subset[j(i)]}e_{J}.$
By the induction hypothesis, the second term is $e_{I_{j(i)}^{*}}$. Thus,
$e_{i}^{*}=\frac{1}{\alpha}\left(\sum_{j=j(i)+1}^{i}\alpha_{j}\right)e_{i}+e_{j(i)}^{*}.$
By construction, $e_{i}^{*}=e_{I_{i}^{*}}$. ∎
Proof of Lemma 4.4.
If $j(i)=0$, the result follows because $e_{i}>e_{j}$ for all $j<i$, $I_{i}^{*}=\{i\}$, and $I_{i-1}^{*}\setminus\{j\in I_{i-1}^{*}:e_{j}>e_{i}\}=\emptyset$. If $j(i)>0$, we are left to show that
$I_{j(i)}^{*}=\{k_{1}(i-1),\ldots,k_{m^{\prime}_{i}}(i-1)\},\quad\text{where }m^{\prime}_{i}=\max\{m:e_{k_{m}(i-1)}\leq e_{i}\}.$ (A.2)
By definition of $j(i)$, for any $j(i)<k<i$, $e_{k}>e_{i}\geq e_{j(i)}$. Thus, for each $j\leq j(i)$,
$\min\{e_{j},\ldots,e_{i-1}\}=\min\{e_{j},\ldots,e_{j(i)}\}.$
(A.2) then follows from the representation (4.2). ∎
A.3 Proofs of results in Section 5
Proof of Lemma 5.2.
The proof is a straightforward manipulation of the expansion of $e_{I}$ using the definition of $w_{i}(I)$, regrouping by the $\alpha_{j}$ terms.
$e_{I}$ $=\frac{1}{\alpha}\sum_{i\in I}\left(\alpha_{i}+\sum_{j\notin I}\alpha_{j}\sum_{p\in\mathcal{P}_{j,i}^{(I)}}\prod_{\ell=1}^{k(p)}q_{i_{\ell-1},i_{\ell}}\right)e_{i}$
$=\frac{1}{\alpha}\sum_{i\in I}\alpha_{i}e_{i}+\frac{1}{\alpha}\sum_{i\in I}\left(\sum_{j\notin I}\alpha_{j}\sum_{p\in\mathcal{P}_{j,i}^{(I)}}\prod_{\ell=1}^{k(p)}q_{i_{\ell-1},i_{\ell}}\right)e_{i}$
$=\frac{1}{\alpha}\sum_{i\in I}\alpha_{i}e_{i}^{(I)}+\frac{1}{\alpha}\sum_{j\notin I}\alpha_{j}\left(\sum_{i\in I}e_{i}\sum_{p\in\mathcal{P}_{j,i}^{(I)}}\prod_{\ell=1}^{k(p)}q_{i_{\ell-1},i_{\ell}}\right)$
$=\frac{1}{\alpha}\sum_{i\in I}\alpha_{i}e_{i}^{(I)}+\sum_{j\notin I}\alpha_{j}e_{j}^{(I)}=\frac{1}{\alpha}\sum_{j=1}^{n}\alpha_{j}e_{j}^{(I)}.$
Proof of Lemma 5.3.
To prove this, we divide the set of paths $\mathcal{P}_{j,i}^{(I)}$ into those with a given first node $i_{1}=k$ which is a child of $j$. These will all have the same first term $q_{j,k}$. Then we know that the resulting set of paths is the same as the set $\mathcal{P}_{k,i}^{(I)}$, but with an additional index $j$ at the start. Then we can rewrite the definition of $e_{j}^{(I)}$ for $j\notin I$ as
$e_{j}^{(I)}$ $=\sum_{i\in I}e_{i}\left(\sum_{p\in\mathcal{P}_{j,i}^{(I)}}\prod_{\ell=1}^{k(p)}q_{i_{\ell-1},i_{\ell}}\right)$
$=\sum_{(j,k)\in E}q_{j,k}\sum_{i\in I}e_{i}\left(\sum_{p\in\mathcal{P}_{j,i}^{(I)},i_{1}=k}\prod_{\ell=2}^{k(p)}q_{i_{\ell-1},i_{\ell}}\right)$
$=\sum_{(j,k)\in E}q_{j,k}\sum_{i\in I}e_{i}\left(\sum_{p\in\mathcal{P}_{k,i}^{(I)}}\prod_{\ell=1}^{k(p)}q_{i_{\ell-1},i_{\ell}}\right)=\sum_{(j,k)\in E}q_{j,k}e_{k}^{(I)}.$
Proof of Theorem 5.4.
We first prove the first part by induction backward from $j=n$ to $j=1$, for any $i\in[n]$, fixing $I\ni i$. For the base case $j=n$, we consider two cases. If $i=n$, it is clear that $e_{i}^{(I)}\geq e_{i}=e_{i}^{(i)}$. If $i<n$, since the hypotheses are topologically ordered, $n\notin A_{i}$ and hence $e_{n}^{(i)}=0\leq e_{n}^{(I)}$. Assuming that $e_{k}^{(I)}\geq e_{k}^{(i)}$ for all $k>j$, and we wish to show that $e_{j}^{(I)}\geq e_{j}^{(i)}$. If $j\notin A_{i}$ we have it automatically as $e_{j}^{(i)}=0$, and otherwise,
$e_{j}^{(i)}=\min\left(e_{j},\sum_{(j,k)\in E}q_{jk}e_{k}^{(i)}\right)\leq e_{j}^{(I)},$ (A.3)
as
$e_{j}^{(I)}=\begin{cases}e_{j}&j\in I\\ \sum_{(j,k)\in E}q_{jk}e_{k}^{(I)}&j\not\in I\end{cases},$
and $e_{j}^{(i)}\leq e_{j}$ and $e_{j}^{(i)}\leq\sum_{(j,k)\in E}q_{jk}e_{k}^{(i)}\leq\sum_{(j,k)\in E}q_{jk}e_{k}^{(I)}$ by the inductive assumption, meaning regardless of whether $j\in I$, $e_{j}^{(i)}\leq e_{j}^{(I)}$ as intended.
For the second part, let
$I_{i}^{(i)}=\{i\},\quad I_{j}^{(i)}=\begin{cases}I_{j+1}^{(i)}\cup\{j\}&e_{j}^{(i)}=e_{j},j\in A_{i}\\ I_{j+1}^{(i)}&\text{otherwise}\end{cases}.$
We then construct $I_{i}^{*}=I_{1}^{(i)}$. For any $j\notin I_{i}^{*}$, there is no path from $j$ to any element in $I$. By Definition 5.1, $e_{j}^{(I_{i}^{*})}=0$ for any $j>i$. We prove $e_{j}^{(i)}=e_{j}^{(I_{i}^{*})}$ for $j\in A_{i}$ by induction backward from $j=i$ to $j=1$, for each $i\in[n]$. For the base case $j=i$, $e_{i}^{(i)}=e_{i}$, which is $e_{i}^{(I_{i}^{*})}$ by Definition 5.1. Assuming that $e_{k}^{(i)}=e_{k}^{(I_{i}^{*})}$ for all $k\in A_{i}$ and $k>j$, we wish to prove $e_{j}^{(i)}=e_{j}^{(I_{i}^{*})}$. If $j\in I_{i}^{*}$, $e_{j}^{(I_{i}^{*})}=e_{j}$ by Definition 5.1 and $e_{j}^{(i)}=e_{j}$ by construction of $I_{i}^{*}$. If $j\notin I_{i}^{*}$,
$e_{j}^{(i)}=\sum_{(j,k)\in E}q_{j,k}e_{k}^{(i)}.$
By the induction hypothesis, $e_{k}^{(i)}=e_{k}^{(I_{i}^{*})}$ for all $k$ such that $(j,k)\in E$. By Lemma 5.3,
$e_{j}^{(i)}=\sum_{(j,k)\in E}q_{j,k}e_{k}^{(I_{i}^{*})}=e_{j}^{(I_{i}^{*})}.$
The proof is then completed.
Appendix B Extensions to select non-DAGs
In this appendix we extend our result from Section 5 to a larger class of graphs which we call Index-Local DAGs. The key motivating observation is that our argument in Section 5 relies on the fact that for every index $i$, the graph over which we have to optimize is a DAG, with $i$ as its only leaf. Thus, we define:
Definition B.1.
A graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ is an Index-Local DAG (ILDAG) if for every $i$, the graph $\mathcal{G}^{(i)}$ formed by removing $j\in V$ such that there is no path from $j$ to $i$, as well as removing any $(i,k)\in\mathcal{E}$, is a DAG.
The motivation for this is that when computing $e_{i}^{*}$, we have shown we only need to consider subsets of the ancestor graph $I\subset A_{i}$, and moreover only subsets containing $i$. Since $i\in I$, we know the $\alpha$-budget at $i$ will never be redistributed, i.e. $e_{i}^{(i)}=e_{i}$ in the notation of Section 5, so we can remove the edges from $i$ and have the same minimization problem. At this point, if the remaining graph is a DAG then we have by the result of Section 5 that the DAG algorithm, for each $i$, will produce correct adjusted e-values $\{e_{i}^{*}\}_{i=1}^{n}$. This can be done by performing the DAG algorithm’s backwards search for each $i$ on its ancestor graph $A_{i}$, and ignoring any reverse edges $i\rightarrow j$.
The result is somewhat limited, but there are at least two interesting cases. The first is the cyclical Fallback graph, which is identical but for an additional edge $n\rightarrow 1$, ensuring that we are always working with the full $\alpha$-budget. This graph clearly satisfies the requirement to be an ILDAG, since removing any one edge $(i\rightarrow i+1)$ breaks the cycle and forms a Fallback chain starting at $i+1$ and ending at $i$.
The other example is the gatekeeper procedure [e.g. Bretz et al., 2009] for two endpoints with a cycle between them, a visualization of which is in Figure 6.
Figure 6: A gatekeeper graph with two nodes in each row. It is an example of an Index-Local DAG satisfying Definition B.1 since at least one of the two cycle-producing edges $H_{3},H_{4}$ is removed when turning any given node into a leaf.
Because we rerun Algorithm 4 for each $i\in[n]$, with potentially all nodes included in $A_{i}$, a naïve upper bound on the worst-case runtime is $O(n^{2}|\mathcal{E}|)$.
Appendix C Expected runtime of e-Fallback algorithm 2
Theorem C.1.
In the context of Algorithm 2:
  1. 1.
    If the e-values $\{e_{i}\}_{i=1}^{n}$ are exchangeable, then the runtime is $O_{P}(n)$.
  2. 2.
    If there are $k$ non-null e-values, and the $n-k$ null e-values are exchangeable, then the runtime is $O_{P}(kn)$.
Proof.
First we show part (1), by denoting $t_{i}$ as the number of comparisons to $e_{i}$ queried, i.e. $t_{i}=i-j(i)$ in the notation of the previous proof. Then we see that the runtime is $O(\sum_{i=1}^{n}t_{i})$, where in expectation we compute
$\mathbb{E}_{P}t_{i}=\sum_{k=1}^{\infty}\mathbb{P}(t_{i}\geq k)=\sum_{k=1}^{i-1}\mathbb{P}(t_{i}\geq k),$ (C.1)
where $\mathbb{P}(t_{i}\geq k)=\mathbb{P}(\min(\{e_{i-k+1},...,e_{i}\}=e_{i})\leq\frac{1}{k!}$ with equality if the distribution of $e_{i}$ is continuous, a result following from order statistics of exchangeable collections of random variables. By the Darth Vader Rule, $\mathbb{E}_{P}[t_{i}]\leq\sum_{k=1}^{i-1}\frac{1}{k!}\leq e-1$, so we conclude that our runtime is $O_{P}((e-1)n)=O_{P}(n)$.
For part (2), we use a similar framework, but lower bound $t_{i}\leq n$ for $e_{i}$ a non-null, and $t_{i}\leq k+t_{i}^{\prime}$ where $t_{i}^{\prime}$ is the number of nulls preceding $i$ until we reach a null e-value with $e_{j}\leq e_{i}$. We enumerate the nulls by indices $\{i_{j}\}_{j=1}^{n-k}$, and define for each $i=i_{\ell}$ $j_{0}:=\max\{j<\ell|e_{i_{j}}\leq e_{i}\}$, the corresponding index for just the nulls. Then defining $t_{i}^{\prime}=\ell-j_{0}$ we have $\mathbb{E}_{P}[t_{i}^{\prime}]\leq e-1$ again, and our upper bound $t_{i}\leq k+e-1$. Putting these together for $k$ non-nulls and $n-k$ nulls, we have runtime
$O_{P}(k(n)+(n-k)(k+e-1))$ $=O_{P}(nk+nk+(e-1)n-k^{2}-(e-1)k)$
$=O_{P}(nk+n),$
so $O_{P}(nk)$ for $k>0$.
This result is nice but uses exchangeability assumptions that may be unrealistic. However, it gives a sense of why we expect that this algorithm will provide a significant runtime improvement, which is achieved explicitly with the modified algorithm in Section 4.3.
Appendix D e-Holm adjusted e-values algorithm
While the procedure 3.1 for the rejection set has an $O(n)$ complexity, it does not compute adjusted e-values. In this appendix, we re-derive and restate the e-Holm algorithm given by Vovk and Wang [2021, 2023]. By (2.5),
$e_{i}^{*}=\min_{I\ni i}\frac{1}{|I|}\sum_{j\in I}e_{j}=\min_{I\ni i}\frac{1}{|I|}\left\{e_{i}+\sum_{j\in I\setminus\{i\}}e_{j}\right\}.$
Clearly, the minimum is achieved by including in $I$ the $|I|-1$ smallest e-values other than $e_{i}$. To pin down the exact expression, we sort our e-values $e_{(1)},...,e_{(n)}$ in decreasing order (to match the order of the associated 1/e p-values), and call $H_{(i)}$ the hypothesis associated with the e-value $e_{(i)}$ and adjusted e-value $e_{(i)}^{*}$. Then
$\min_{I\ni i,|I|=k+1}\frac{1}{|I|}\sum_{j\in I}e_{(j)}=\begin{cases}\frac{e_{(i)}+\sum_{j=n-k+1}^{n}e_{(j)}}{k+1}&k\leq n-i\\ \frac{\sum_{j=n-k}^{n}e_{(j)}}{k+1}&k>n-i\end{cases}.$
To simplify the notation, let
$E_{k}=\sum_{j=n-k+1}^{n}e_{(j)}.$
Then we can express the adjusted e-value for $H_{(i)}$ as
$e_{(i)}^{*}=\min\left\{\min_{k\leq n-i}\frac{e_{(i)}+E_{k}}{k+1},\min_{k>n-i}\frac{E_{k+1}}{k+1}\right\}=\min_{k\leq n-i}\frac{e_{(i)}+E_{k}}{k+1},$ (D.1)
where the second equality follows because $E_{k}/k$ is non-decreasing and thus, for any $k>n-i$,
$\min_{k>n-i}\frac{E_{k+1}}{k+1}\geq\frac{E_{n-i+1}}{n-i+1}=\frac{e_{(i)}+E_{n-i}}{n-i+1}\geq\min_{k\leq n-i}\frac{e_{(i)}+E_{k}}{k+1}.$
The terms $E_{1},\ldots,E_{n}$ can be jointly computed in $O(n\log n)$ time, including sorting and cumulative sums. Calculating all adjusted e-values separately would result in $O(n^{2})$ time in total because $e_{(i)}^{*}$ is the minimum of $n-i$ terms. We can reduce the computation cost to $O(n)$ by leveraging the following observation.
Lemma D.1.
Let $k_{i}=\max\{k:e_{(i)}^{*}=(e_{(i)}+E_{k})/(k+1)\}$. Then $k_{i}$ is non-increasing in $i$.
Proof.
By definition of $k_{i}$, for any $k>k_{i}$,
$\frac{e_{(i)}+E_{k}}{k+1}>\frac{e_{(i)}+E_{k_{i}}}{k_{i}+1}\Longrightarrow\left(\frac{1}{k_{i}+1}-\frac{1}{k+1}\right)e_{(i)}<\frac{E_{k}}{k+1}-\frac{E_{k_{i}}}{k_{i}+1}.$
Since $e_{(i)}\geq e_{(i+1)}$,
$\left(\frac{1}{k_{i}+1}-\frac{1}{k+1}\right)e_{(i+1)}<\frac{E_{k}}{k+1}-\frac{E_{k_{i}}}{k_{i}+1}\Longrightarrow\frac{e_{(i+1)}+E_{k}}{k+1}>\frac{e_{(i+1)}+E_{k_{i}}}{k_{i}+1}.$
This implies the minimum of $(e_{(i+1)}+E_{k})/(k+1)$ cannot be attained for some $k>k_{i}$. As a result, $k_{i+1}\leq k_{i}$. ∎
Lemma D.1 suggests $k_{i}$ and the adjusted e-values can be computed sequentially in a decreasing order from $i=n$ to $1$. The details are presented in Algorithm 5.
This algorithm correctly computes our minima as earlier argued, and has runtime $O(n\log n)$, from the sorting. The computation only changes $k$’s value $n$ times, and changes $e^{*}$ a maximum of $2n$ times.
Input: Vectors $\{e_{i}\}_{i=1}^{n}$;
Sort: $\{e_{i}\}_{i=1}^{n}$ from highest to lowest as $\{e_{(i)}\}_{i=1}^{n}$;
Compute: $E_{1}=e_{(n)}$, $E_{k}=E_{k-1}+e_{(n-k+1)}$ for $k=2$ to $k=n$;
Initialize: $e_{(n)}^{*}=e_{(n)}$, $e^{*}=e_{(n)},k=1$;
For: $i=n-1$ to 1,
Do: $e^{*}=e^{*}+\frac{e_{(i)}-e_{(i+1)}}{1+k}$;
While ($e^{*}>e_{(n-k)}$): $e^{*}=\frac{1+k}{2+k}e^{*}+\frac{1}{2+k}e_{(n-k)}$; $k=k+1$.
$e_{(i)}^{*}=e^{*}$.
Output: $\{e_{(i)}^{*}\}_{j=i}^{n}$.
Algorithm 5 e-Holm Adjusted e-values